Is the function f : R ➡R, defined by f(x) = 1-|x| is differentiable at x=1

1
Expert's answer
2021-11-02T18:36:27-0400
"\lim\limits_{h\to0^-}\dfrac{f(1+h)-f(1)}{h}=\lim\limits_{h\to0^-}\dfrac{1-|1+h|-1+|1|}{h}"

"=\lim\limits_{h\to0^-}\dfrac{-h}{h}=-1"



"\lim\limits_{h\to0^+}\dfrac{f(1+h)-f(1)}{h}=\lim\limits_{h\to0^+}\dfrac{1-|1+h|-1+|1|}{h}"

"=\lim\limits_{h\to0^+}\dfrac{-h}{h}=-1"

"\lim\limits_{h\to0^-}\dfrac{f(1+h)-f(1)}{h}=-1=\lim\limits_{h\to0^+}\dfrac{f(1+h)-f(1)}{h}"

Then


"\lim\limits_{h\to0}\dfrac{f(1+h)-f(1)}{h}=-1"

Therefore the function "f(x) = 1-|x|" is differentiable at "x=1" and "f'(1)=-1."